OpenStudy (anonymous):

If B is between A and C, AB=8 and AC=21, what is the value of BC?

OpenStudy (anonymous):

make another equation ... B-A=X, C-B=x, C-A = 2X

OpenStudy (anonymous):

8c/b=21

OpenStudy (anonymous):

13